### v1

# LEO-VL: Towards 3D Vision-Language Generalists via Data Scaling with Efficient Representation

Abstract
Developing 3D-VL generalists capable of understanding 3D scenes and following natural language instructions to perform a wide range of tasks has been a longstanding goal in the 3D-VL community. Despite recent progress, 3D-VL models still lag behind their 2D counterparts in capability and robustness, falling short of the generalist standard. A key obstacle to developing 3D-VL generalists lies in data scalability, hindered by the lack of an efficient scene representation. We propose LEO-VL, a 3D-VL model built upon condensed feature grid (CFG), an efficient scene representation that bridges 2D perception and 3D spatial structure while significantly reducing token overhead. This efficiency unlocks large-scale training towards 3D-VL generalist, for which we curate over 700k high-quality 3D-VL data spanning four domains of real-world indoor scenes and five tasks such as captioning and dialogue. LEO-VL achieves state-of-the-art performance on a variety of 3D QA benchmarks, including SQA3D, MSQA, and Beacon3D. Ablation studies confirm the efficiency of our representation, the importance of task and scene diversity, and the validity of our data curation principle. Furthermore, we introduce SceneDPO, a novel post-training objective that enhances the robustness of 3D-VL models. We hope our findings contribute to the advancement of scalable and robust 3D-VL generalists.
![[Uncaptioned image]](x1.png)
1 Introduction
With the crucial role of 3D scene understanding in human intelligence, building generalist models that can understand 3D environments and communicate through natural language [34, 97, 78, 55] has become a central goal of the 3D vision-language (3D-VL) community. Recent progress has advanced capabilities across 3D-VL tasks such as grounding [117, 118, 39], captioning [53, 93], reasoning [32, 46], and dialogue [34, 95]. Nonetheless, current 3D-VL models fall significantly short of the generalist capabilities demonstrated by their 2D counterparts [60, 82, 88, 13, 42]. We posit that a key obstacle to developing 3D-VL generalists lies in data scalability, hindered by the absence of an efficient scene representation. Therefore, establishing an efficient and scalable scene representation is a critical first step towards building 3D-VL generalists.
Existing scene representations in 3D-VL models can be categorized into two approaches. The first adopts 3D modality inputs such as point clouds, but requires complex pre-processing pipelines such as 3D reconstruction and instance segmentation [117, 39, 11]. And they confront the inherent challenge of direct 3D perception, which is exacerbated by the scarcity of 3D-VL data. The other approach leverages 2D vision-language (2D-VL) models for perception [115, 111, 71], fueled by the advancement of 2D perception. However, this approach suffers from poor data scalability, as representing a 3D scene with multi-view 2D features incurs significant computational overhead. Additionally, both approaches typically use rudimentary 3D position embeddings for 3D spatial modeling, which fall short in capturing 3D spatial structures effectively.
We propose LEO-VL, a 3D-VL model built on condensed feature grid (CFG), an efficient scene representation that bridges 2D perception and 3D spatial structure, significantly improving efficiency while preserving the strength of 2D perception. Specifically, we employ a pretrained 2D visual encoder to extract image features, which are then back-projected into 3D space for voxelization. We utilize rotary position embedding (RoPE) [79] to encode the height of voxels and condense voxel features vertically to form a CFG. For horizontal spatial modeling, we integrate 2D Fourier features [80] into the CFG. In addition to efficiency, our CFG simplifies 3D spatial modeling by disentangling vertical and horizontal dimensions. The scene tokens of CFG are further connected to a large language model (LLM) for auto-regressive language modeling.
For the training of LEO-VL, we incorporate five 3D-VL tasks, including captioning, question answering (QA), planning, and dialogue, spanning four domains of real-world indoor scenes: ScanNet [19], 3RScan [86], MultiScan [58], and ARKitScenes [5]. Within each domain, we curate existing datasets and newly generated data while prioritizing data quality over scale. This could avoid model performance degradation from naive scaling with low-quality sources. With our curation principle, we collect a balanced and diverse corpus of over 700k high-quality 3D-VL data samples for the instruction tuning of LEO-VL.
We evaluate LEO-VL on a variety of 3D QA benchmarks, demonstrating state-of-the-art performance with significantly higher efficiency than prior methods. Ablation studies exhibit advantages of the CFG in efficiency and 3D spatial modeling. Ablations on various training data configurations reveal the benefits of tasks such as captioning and dialogue, and the importance of scene diversity for strong cross-domain performance. We further show that scaling with simplistic QA data leads to performance degradation, confirming our data curation principle, while our curated data yields consistent scaling effects. Additionally, we introduce SceneDPO, a novel post-training objective that incorporates contrasts across answers and scenes, and demonstrate its effectiveness in enhancing model robustness compared to conventional supervised finetuning (SFT).
In summary, our contributions are as follows:
1.
We propose LEO-VL, a 3D-VL generalist built upon condensed feature grid (CFG), an efficient scene representation that enables strong capability in 3D scene perception with significantly improved efficiency and simplified 3D spatial modeling.
2.
We construct a comprehensive 3D-VL instruction-tuning dataset spanning four scene domains and five tasks. Our dataset comprises over 700k high-quality samples, striking a balance between scale and quality, and serving as a reliable foundation for the training of 3D-VL generalists.
3.
We demonstrate LEO-VL’s state-of-the-art performance across 3D QA benchmarks such as SQA3D. Ablation studies validate the efficiency of our representation, the benefits of task and domain diversity, and confirm our data curation principle. We also showcase the superiority of our post-training objective in enhancing model robustness compared to conventional SFT.
2 Related Work
3D Vision-Language Models.
Early work in 3D-VL understanding develops models that process 3D representations such as point clouds [67, 69, 65, 45, 91, 68, 108, 37, 100] and voxels [27, 59, 74, 81, 28, 18, 15, 76]. With the maturation of pretraining and instruction tuning techniques, 3D-VL models have demonstrated significant performance improvements [109, 36, 1, 10, 36, 49, 94, 116, 117, 114, 90] and the capability of generalist models [29, 93, 34, 95, 11, 70, 118, 26, 107, 16, 32, 21]. Nonetheless, 3D perception remains challenging due to the data scarcity and inherent complexity of 3D representation space. In contrast, an alternative line of work leverages 2D perception models to handle 3D-VL tasks [64, 77, 23, 38, 31, 24, 57, 52], a strategy made increasingly effective by advances in 2D vision-language models [48, 88, 106, 115, 111, 71]. However, adapting 2D VLMs for 3D scene understanding poses two challenges: the cost of representing 3D scenes and the absence of effective 3D spatial modeling. While recent efforts [113, 99, 83, 33] attempt to address these issues, they still fall short of achieving both high efficiency and strong capability. In this work, we propose a framework that balances efficiency and effectiveness, showing the potential of foundational 3D-VL generalist models.
3D Vision-Language Datasets and Benchmarks.
3D-VL datasets and benchmarks are typically built upon 3D scene assets [19, 98, 86, 58, 5, 7, 73, 8, 41, 112, 25, 20]. Existing 3D-VL datasets primarily focus on object grounding [9, 2, 105] and QA [3, 56] tasks, but most of them are centered around ScanNet [19]. As the 3D-VL community advances, recent efforts have focused on aggregating diverse scene domains to enable large-scale pretraining [117, 118, 39, 89] and generalist instruction tuning [34]. This further drives the development of cross-domain datasets [96, 54, 46, 78, 102] and benchmarks [46, 35, 97]. Following this trend, our simple yet efficient framework is the first to scale 3D-VL generalist learning to the SceneVerse level [39], covering four scene domains [19, 86, 58, 5] and five 3D-VL tasks. We further demonstrate its effectiveness through comprehensive evaluations across these heterogeneous 3D scenes.
Vision-Language Learning Objective.
Prior studies [22, 35, 63, 104] indicate that SFT, as a common learning objective for 3D-VL models, may undergo weak robustness given the scarce data and overfitting risk in 3D-VL learning. Existing efforts to remedy the robustness mainly involve data augmentation [96, 40, 110], with the learning objective underexplored. Motivated by the success of reinforcement learning from human feedback (RLHF) [61, 75, 72, 101, 62, 50, 17] for post-training, prior 2D-VL models [43, 66, 87, 92] employ direct preference optimization (DPO) objective [72] to address the robustness issue. We bring these insights into 3D-VL learning and introduce an additional objective with scene contrast to post-training for more robust 3D-VL models.
3 Method
In this section, we outline our approach to 3D-VL generalists. In Section 3.1, we introduce our model, LEO-VL, and particularly its efficient representation of 3D scenes, which largely reduces the token overhead while preserving strong perception capability. In Section 3.2, we introduce the recipe of our comprehensive instruction-tuning datasets for training of 3D-VL generalists. In Section 3.3, we further introduce a novel post-training objective to enhance the robustness of 3D-VL models.
3.1 Model
As illustrated in Fig. 2, LEO-VL extracts image features from RGBD frames and transforms the features into an efficient scene representation, termed the condensed feature grid (CFG). Next, an LLM performs auto-regressive language modeling over the joint sequence of CFG tokens and instruction tokens. In contrast to prior 3D-VL models that rely on object-centric 3D representation [34, 118, 32, 46], LEO-VL adopts 2D perception considering the limitations of the object-centric approach, such as the dependency of instance segmentation and loss of scene-level information.
2D Perception and Back-projection.
We employ a 2D visual encoder to extract image features from multi-view RGB inputs. Given a set of RGB images , the encoder produces corresponding feature maps , where , , and denote the resolution and feature dimension. For each view, the corresponding depth maps are used to back-project 2D pixels into 3D space. Let and denote the intrinsic and extrinsic camera matrices, respectively. A pixel at coordinate with depth is back-projected to a 3D point through the following homogeneous transformation:
We compute 3D coordinates for all pixels and interpolate them to align with the feature map resolution . These 3D points are then associated with corresponding 2D features.
Condensed Feature Grid.
First, we voxelize the 3D points and average pool the point features within each voxel, discarding empty voxels. For referring tasks such as object captioning, a learnable anchor embedding is injected into voxels that fall within the referred region. Despite voxelization, the representation remains costly and still limits scalability. To further improve efficiency, we convert voxels into the CFG by pooling voxel features along the height axis. Before pooling, we encode the height of each voxel with RoPE [79]. To encode horizontal positions, we integrate 2D Fourier features [80] in the CFG.
Position Embeddings.
We adopt RoPE for vertical spatial modeling and Fourier features for horizontal spatial modeling. RoPE encodes position information by rotating the feature vector by a series of angles for different feature channels. The angles are derived from linear combinations of the position value and channel-specific frequencies. The rotary nature enables RoPE to better capture vertical spatial relations, such as distinguishing objects at different heights, compared to additive position embeddings [85, 80]. In contrast, the Fourier feature works by first applying a linear transformation to the position value and then computing sinusoidal components, which are passed through an MLP to produce embeddings. Let denote the rotation matrix for RoPE and the projection matrix for the Fourier feature. The formulations of position-encoded features are as follows:
Large Language Model.
3.2 Data
The construction of LEO-VL training data is guided by three principles: (1) the diversity in scene domains, which demonstrates crucial for 3D-VL learning [39]; (2) the diversity of tasks, which we think necessary given the benefits of long-response tasks [34, 54]; and (3) the balance between quality and scale, considering the potential risks in hindering model performance from low-quality data. We detail the composition of LEO-VL training data in Table 1.
Scene Domains.
We include four domains of real-world indoor scenes: ScanNet [19], 3RScan [86], MultiScan [58], and ARKitScenes [5], following prior practices in scaling 3D scenes for 3D-VL learning [39, 46]. We exclude scene domains that lack attribute-rich scene graphs (e.g., HM3D [73]), which are essential for LLM-assisted data generation. We follow the official splits for ScanNet and 3RScan, SceneVerse splits [39] for MultiScan, and MSQA splits [46] for ARKitScenes.
ScanNet | 3RScan | MultiScan | ARKitScenes | Overall count | Avg. length (str) | Avg. length (words) | |
ObjCap | SV, MM | LEO, SV, MM | SV | SV | 216k | 299 | 56 |
SceneCap | MM, ✓ | LEO, MM, ✓ | ✓ | ✓ | 128k | 633 | 104 |
QA | ScanQA, SQA, MSQA | MSQA | - | MSQA | 289k | 31 | 6 |
Plan | ✓ | LEO | ✓ | ✓ | 18k | 534 | 96 |
Dialog | ✓ | LEO | ✓ | ✓ | 61k | 93 | 18 |
Tasks and Datasets.
We incorporate five prevalent 3D-VL tasks with natural language outputs, all compatible with a unified instruction tuning framework. We exclude the 3D object grounding task due to its different formulation. We think the absence of the grounding task is not detrimental given the success of 2D-VL generalists without explicit grounding supervision [47, 60, 88, 13]. In addition to existing datasets, we also generate new data by prompting LLMs with scene graphs [34, 46]. Our data covers five categories of text-output tasks:
Object Captioning. This task requires the model to describe a specific object in natural language. We adopt object captions from SceneVerse [39] and MMScan [54]. We exclude datasets that repurpose object grounding texts as captions [14, 2], as such texts lack diverse descriptions regarding object attributes.
Planning. Given a high-level goal (e.g., “organize the room”), the model is required to generate a step-by-step grounded plan. We use the planning data from LEO [34] for 3RScan, and generate new data for other scene domains.
Dialogue. This task requires the model to generate context-aware responses conditioned on both the 3D scene and dialogue history. We use the dialogue data from LEO [34] for 3RScan, and generate new data for other scene domains.
3.3 Post-training
Comprehensive instruction tuning alone may be insufficient to ensure model robustness, as evidenced by prior studies in 2D-VL models [43, 66, 87]. Given the pronounced risk of overfitting in 3D-VL models [22, 35], it is essential to design an effective post-training objective to enhance their robustness.
To this end, we propose SceneDPO, a novel learning objective for the post-training of 3D-VL models. First, we follow DPO to adopt a loss term that contrasts between positive answer and negative answer . Motivated by the issue of visual ignorance [35] in 3D-VL models, where answers are predicted by overfitting to text without utilizing scene information, we introduce a loss term that contrasts between positive scene and negative scene . This discourages model from predicting current answer when conditioned on irrelevant scenes. To mitigate degradation of positive answers, we also incorporate a negative log-likelihood loss term , which proves critical in prior works [50, 62, 87]. Given a training tuple , the overall loss is defined as follows:
where denotes the trained model, ref denotes the reference model, denotes the sigmoid function, and , , , and are scalar hyperparameters.
4 Experiment
In this section, we first compare LEO-VL against state-of-the-art models on a variety of 3D QA benchmarks in Section 4.1, highlighting its advantages in both efficiency and accuracy. In Section 4.2, we present ablation studies on model design to show the effectiveness of CFG. In Section 4.3, we analyze various training data configurations to reveal the influence of task diversity, scene domain coverage, data quality, and data scale on 3D-VL generalist learning. In Section 4.4, we further explore the effect of post-training objective based on LEO-VL.
Implementation Details.
We initialize the 2D visual encoder and LLM with Qwen2.5-VL-7B-Instruct [4]. The learnable parameters include position embedding, anchor embedding, and LoRA parameters [30] of the LLM, which amount to 66M in total. We set the voxel size to 0.2 meters and retain up to 750 scene tokens of CFG. We train LEO-VL on our comprehensive 3D-VL dataset for 5 epochs, which takes 2 days with 8 NVIDIA A100 80G GPUs. We adopt AdamW optimizer [51] with a base learning rate at , scheduled with linear warmup and cosine decay. We provide detailed hyperparameters in Appendix.
Evaluation.
We primarily evaluate LEO-VL on 3D QA benchmarks considering the limited availability of standardized benchmarks for captioning, planning, and dialogue. Existing object captioning benchmarks [14, 2] are excluded in the main paper (see results in Appendix C), as their captions are derived from object grounding texts with simplistic patterns that cannot reflect natural descriptive capabilities. Our evaluation includes ScanQA [3] for general QA in 3D scenes, SQA3D [56] for situated QA, and MSQA [46] for situated QA across multiple scene domains (i.e., ScanNet, 3RScan, and ARKitScenes). We also incorporate Beacon3D [35], which offers robust metrics for 3D QA evaluation. Evaluation metrics include n-gram metrics (CIDEr and BLEU-4) for ScanQA, exact-match accuracy (EM and EM-R) for SQA3D, and GPT-Score for MSQA and Beacon3D.
4.1 Comparison with State-of-the-art Models
Model | Scene (#tokens) | ScanQA (val) | SQA3D (test) | MSQA (test) | Beacon3D | MSQA (test) | MSQA (test) | |||
C | B-4 | EM | EM-R | GPT-Score | Case | Obj. | GPT-Score | GPT-Score | ||
ScanQA [3] | Query (256) | 64.9 | 10.1 | 47.2 | - | - | - | - | - | - |
3D-LLM [29] | Query (32) | 74.5 | 12.9 | 49.8 | - | - | - | - | - | - |
PQ3D [118] | Query (80) | - | - | 47.1 | - | - | 35.9 | 4.2 | - | - |
DSPNet [52] | Query (256) | - | - | 50.4 | - | - | - | - | - | - |
3D-LLaVA [21] | Query (100) | 92.6 | 17.1 | 54.5 | 56.6 | - | - | - | - | - |
3D-VisTA [117] | Object (80) | 69.6 | 10.4 | 48.5 | - | - | 35.3† | 8.1† | - | - |
LEO [34] | Object (60) | 101.4 | 13.2 | 50.0 | 52.4 | 56.7† | 45.2 | 7.5 | 50.4† | 59.7† |
SceneVerse [39] | Object (80) | - | - | 49.9 | - | - | 40.5 | 4.7 | - | - |
Chat-Scene [32] | Object (200) | 87.7 | 14.3 | 54.6 | 57.5 | - | 45.8† | 7.8† | - | - |
Inst3D-LMM [99] | Object (200) | 88.6 | 14.9 | - | - | - | - | - | - | - |
Scene-LLM [26] | Voxel (0.18m) | 80.0 | 11.7 | 53.6 | - | - | - | - | - | - |
LLaVA-3D [115] | Voxel (3096) | 91.7 | 14.5 | 55.6 | 57.6 | - | - | - | - | - |
Video-3D LLM [111] | Video (32 frames) | 100.5 | 16.3 | 57.7 | - | - | - | - | - | - |
GPT4Scene [71] | Video (32 frames) | 96.3 | 15.5 | 59.4 | 62.4 | - | - | - | - | - |
LEO-VL | Grid (750) | 97.6 | 14.8 | 59.7 | 62.6 | 58.2 | 59.5 | 18.2 | 51.4 | 64.2 |
Baselines.
For comparison, we include state-of-the-art 3D-VL models across four categories: (1) query-based methods, ranging from early work like ScanQA [3] to recent 3D-LLaVA [21]; (2) object-centric methods, from 3D-VisTA [117] to Inst3D-LMM [99]; (3) voxel-based methods, such as Scene-LLM [26] and LLaVA-3D [115]; and (4) video-based methods, including Video-3D LLM [111] and GPT4Scene [71]. We also report each model’s scene representation and its associated cost. For Scene-LLM, only the voxel size (0.18 meters) is available, which we estimate to yield a similar number of tokens as LLaVA-3D (voxel size 0.2 meters). For video-based methods, we list the number of frames, noting that 32 frames typically amount to over 6k scene tokens.
Results.
As shown in Section 4.1, LEO-VL achieves state-of-the-art performance on most 3D-VL benchmarks with a low representation cost. While video-based models show competitive performance, their representations involve thousands of tokens and limit data scalability, hindering the acquisition of generalist capabilities through large-scale training. In contrast, the efficient representation of LEO-VL enables 3D-VL data scaling for consistently strong performance across domains including ScanNet, 3RScan, and ARKitScenes. These results underscore the advantages of our efficient representation in perception capability and data scalability.
4.2 Model Analysis
Statistics.
In Fig. 4, we present statistics of voxel tokens and CFG tokens on ScanNet. In addition to the distribution of token counts, we report two metrics: compression rate, defined as the ratio of CFG tokens to raw voxel tokens before condensation; and preservation rate, defined as the proportion of voxels retained by the CFG tokens. The results show that CFG achieves an average compression rate of about and an average preservation rate close to 99%. This demonstrates the high efficiency of our CFG representation.
Efficiency vs. Accuracy.
We provide a joint visualization of efficiency and accuracy in Fig. 3. Accuracy is measured using the EM on SQA3D, the metric with the most available reference results. As shown in Fig. 3, LEO-VL achieves a Pareto optimum between efficiency and accuracy. To investigate the efficiency of CFG, we ablate the representation with two alternatives on the ScanNet subset: voxel-full, which retains all voxel tokens without condensation; and voxel-sample, which downsamples voxel tokens to match the token count of CFG. The results in Fig. 3 show that compared to voxel-full, voxel-sample suffers a drop in accuracy, while CFG even slightly improves the accuracy. These results suggest that high accuracy in 3D-VL tasks can be achieved with significantly lower token costs, as enabled by our efficient CFG representation.
3D Spatial Modeling.
A key advantage of LEO-VL lies in the disentanglement of 3D spatial modeling across vertical and horizontal dimensions. To assess the effectiveness of position embeddings, we train LEO-VL on the ScanNet subset under two ablated settings: “w/o z-pos”, where the z-axis position embedding is removed; and “w/o xyz-pos”, where both z-axis and xy-plane position embeddings are removed. As shown in Fig. 5, both types of position embeddings play critical roles in 3D-VL tasks. Removing the z-axis position embedding leads to a general performance drop, while removing the xy-plane position embedding especially degrades performance on SQA3D. These results validate the effectiveness of our design in 3D spatial modeling.
4.3 Data Analysis
We conduct ablative studies on data to answer the following questions: (1) What is the effect of captioning, planning and dialogue tasks in 3D-VL generalist learning? (2) What is the benefit of scaling across diverse scene domains? (3) What happens when we naively scale up 3D-VL data without considering data quality? (4) How effective is data scaling for LEO-VL with curated data?
Task Effects.
Based on LEO-VL (ScanNet), we ablate the training data with different task configurations: “w/o cap” removes both object captioning and scene captioning tasks, while “w/o pl+dl” removes planning and dialogue tasks. As shown in Fig. 6, both task categories contribute meaningfully to the generalist capabilities of LEO-VL. Captioning proves particularly beneficial for SQA3D, whereas planning and dialogue tasks are more influential for ScanQA. We further perform a qualitative test by prompting “w/o cap” model with a scene captioning task. The model struggles to generate meaningful scene descriptions, likely due to the lack of training in captioning. These findings underscore the importance of task diversity for both acquiring task-specific skills and enhancing cross-task performance, thereby playing a critical role in achieving generalist capabilities.
Data | ScanQA | SQA3D | MSQA | Beacon3D | MSQA | MSQA | Avg. |
ScanNet | 58.7 | 60.3 | 58.4 | 37.9 | 46.1 | 54.7 | 52.7 |
Scan+3R | 56.5 | 59.8 | 56.0 | 37.5 | 51.0 | 56.1 | 52.8 |
All | 56.2 | 61.2 | 58.2 | 38.9 | 51.4 | 64.2 | 55.0 |
Scan+3R | 56.5 | 59.8 | 56.0 | 37.5 | 51.0 | 56.1 | 52.8 |
+QA | 55.7 | 58.3 | 56.8 | 37.5 | 51.3 | 55.6 | 52.5 |
Count | Top-15 coverage | |
Scan+3R | 253k | 12.8% |
Extra QA | 669k | 39.2% |
Scene Domains.
To assess the impact of scaling across scene domains, we introduce an additional checkpoint trained on ScanNet and 3RScan (Scan+3R), complementing existing LEO-VL and LEO-VL (ScanNet). As shown in Section 4.3, compared to “ScanNet”, “Scan+3R” yields significant improvements on 3RScan and ARKitScenes, with only a slight performance drop on ScanNet. Further, compared to “Scan+3R”, “All” shows substantial gains on ARKitScenes. These results suggest that the performance gains may primarily stem from exposure to corresponding domains, emphasizing the importance of a wide domain coverage for robust cross-domain generalist performances.
Priority of Quality over Scale.
Based on the results of “Scan+3R”, we explore the outcome of naively scaling up 3D-VL data without considering data quality. Specifically, we augment our default subsets on ScanNet and 3RScan with extra 3D QA data from two sources: (1) LEO [34], which provides 94k QA samples from 3RScan; and (2) MMScan [54], which offers 575k QA samples from ScanNet and 3RScan. The resulting checkpoint is denoted as “+QA”. Notably, these two datasets exhibit a highly repetitive and simplistic answer distribution. To quantify this, we normalize the answers by replacing common entities with placeholders (e.g., [COLOR]), transforming them into answer templates. As shown in Section 4.3, 39.2% of answers in “+QA” fall into the top-15 frequent templates, compared to only 12.8% in our curated data, demonstrating the simplistic distribution of “+QA”. The results in Section 4.3 reveal that despite incorporating substantially more QA data, “+QA” shows degradation in the overall performance compared to “Scan+3R”. This confirms that scaling with low-quality data can be detrimental, reinforcing our core curation principle: prioritizing data quality over sheer scale is essential for effective training of 3D-VL models.
Consistent Scaling Effects.
To probe the scaling effects, we ablate the proportions of training data based on full-domain coverage, specifically training with 12.5%, 25%, 50%, and 100% of the dataset. We report the overall performance by averaging metrics across all evaluated benchmarks. As shown in Fig. 7, the results exhibit a steady upward trend, demonstrating consistent improvements as the data scale increases. This highlights a desirable scaling behavior enabled by our data design, echoing the noisy trend or even degradation in performance when scaling with low-quality 3D-VL data. The consistent scaling effects underscore the importance of data curation in scaling for 3D-VL generalist training. Furthermore, the continued upward trend with no signs of saturation suggests that further performance gains may be achievable by scaling up more curated 3D-VL data.
4.4 Post-training
Settings.
We explore the effect of post-training objective based on LEO-VL after instruction tuning. Specifically, we continue to fine-tune LEO-VL on SQA3D with SFT and SceneDPO, respectively. To construct negative answers , we first let LEO-VL generate predictions on the training set of SQA3D, and then employ GPT-4o [60] to rephrase correct predictions into hard negative samples. These rephrased negatives, along with the original incorrect predictions, form the negative answer set for SceneDPO. For negative scenes , we randomly sample a different scene from the training set. We evaluate the performance by two metrics: in-domain performance, computed as the average of EM and EM-R on SQA3D; and overall performance, averaged across all benchmarks. We remove the reference model in our implementation, given the empirical observation that it has a negligible impact while largely reducing resource costs. We set , and train both SFT and SceneDPO for 10 epochs.
Data | SQA3D | Avg. |
LEO-VL | 61.2 | 55.0 |
SFT | 62.5 | 53.7 |
SceneDPO | 62.7 | 53.8 |
Results.
The results are reported in Section 4.4 and Fig. 8. As shown in Section 4.4, SceneDPO improves both the in-domain performance on SQA3D and the overall performance across benchmarks, demonstrating its advantage in enhancing post-training performance. Fig. 8 shows consistent improvements in the accuracy over both answer and scene, indicating the acquisition of discriminative ability for the model. Notably, the significant gain in the accuracy over scene suggests the model’s evolution in addressing the issue of visual ignorance. On the other hand, the initially low accuracy also reveals the model’s limited ability in distinguishing semantically similar answers and correctly leveraging scene information for QA. This observation reinforces the need for effective post-training strategies like SceneDPO in enhancing the robustness of 3D-VL models.
5 Conclusion
We propose LEO-VL, an advanced 3D-VL generalist that excels at general 3D-VL tasks with natural language outputs. LEO-VL is equipped with an efficient scene representation that bridges 2D perception and 3D spatial structure, featuring strong perception capability, high representation efficiency, and simplified 3D spatial modeling. For the generalist learning of LEO-VL, we curate a comprehensive 3D-VL dataset that spans four domains of real-world indoor scenes (i.e., ScanNet, 3RScan, MultiScan, and ARKitScenes) and five tasks, covering captioning, QA, planning, and dialogue. We demonstrate the state-of-the-art performance of LEO-VL on various 3D QA benchmarks such as SQA3D. Ablative analyses reveal the importance of diversity in tasks and scene domains. The exploration with additional simplistic QA data confirms our principle that prioritizes data quality over scale, which naturally leads to consistent scaling effects. We hope our experience and findings in this work can benefit future research in developing 3D-VL generalists. Furthermore, we demonstrate the necessity and advantage of post-training objectives in enhancing the robustness of 3D-VL models.
Appendix A Discussion
Limitations.
Despite strong performance and efficiency, LEO-VL exhibits several limitations. First, the use of the CFG effectively reduces token overhead but may lead to the loss of fine-grained details due to feature pooling, potentially limiting performance on tasks requiring precise understanding. Second, while the CFG’s effectiveness is demonstrated on indoor scenes, its generalization to outdoor or dynamic environments remains underexplored. Third, although our data scaling spans diverse real-world scene domains, the potential of leveraging abundant synthetic scenes is not explored yet.
Broader Impact.
This work introduces a 3D-VL generalist model for 3D scene understanding tasks with natural language outputs, offering a step forward in enabling intelligent systems to comprehend and interact with the physical world. The proposed approach holds promise for applications such as assistive robotics, AR/VR interfaces, and digital twin systems for smart environments. Meanwhile, we recognize potential concerns related to data privacy, computational resource consumption, and misuse in surveillance. We encourage the community to pair technical innovation with ethical foresight. And we will release our code and data to support transparent, reproducible, and responsible research in 3D-VL generalists.
Appendix B Implementation Details
Input Sequence.
Each input to the LLM begins with a system prompt, which is followed by the Scene Tokens, derived from the CFG, and then the Instruction, i.e., the textual input. For situated tasks, the situation description is included within the Instruction, followed by task-specific text such as a question. The model then generates the Response, triggered by the generation prompt “assistant”.
LLM Hyperparameters.
We finetune the LLM using LoRA [30] for parameter-efficient learning, with a rank of 16, scaling factor , and a dropout rate of 0, following LEO [34]. LoRA adapters are applied to all projection matrices: in the attention layers and in the MLP blocks. For inference, we employ beam search with repetition and length penalties. Full hyperparameter settings are provided in Table 7.
Training Hyperparameters.
We present the detailed settings in Table 7. The instruction tuning of LEO-VL is conducted for 5 epochs, and the post-training is conducted for 10 epochs.
GPT-Score.
We calculate GPT-Score using the Azure OpenAI API with GPT-4o-2024-08-06 as the evaluator. Each score is evaluated based on a triplet consisting of a question, a ground-truth answer, and a predicted answer. We adopt the scoring prompt from MSR3D [46], which includes a system instruction and example contexts to guide the evaluation criteria.
Appendix C Additional Results
Scan2Cap.
Scan2Cap [14] is a benchmark for 3D object captioning, which utilizes the referential text from ScanRefer [9] as captions. We have not included Scan2Cap results in the main paper as the captions are limited in linguistic diversity and offer limited insight into a model’s descriptive capability. Nonetheless, for completeness, we report the performance of LEO-VL finetuned on Scan2Cap in Appendix C. The results show that LEO-VL achieves state-of-the-art performance on Scan2Cap, attaining the highest scores on two out of four evaluation metrics, highlighting its strong capability in object captioning.
Hyperparameter | Value |
Training strategy | LoRA |
Rank | 16 |
Alpha | 16 |
Dropout | 0.0 |
Inference strategy | Beam search |
Number of beams | 5 |
Max new tokens | 256 |
Repetition penalty | 3.0 |
Length penalty | 1.0 |
Hyperparameter | Value |
Optimizer | AdamW |
Weight decay | 0.05 |
Betas | [0.9, 0.999] |
Base learning rate | |
Warmup steps | 400 |
Type of GPUs | NVIDIA A100/A800 80G |
Number of GPUs | 8 |
Batch size per GPU (total) | 2 (16) |
Gradient accumulation steps | 5 |
Training precision | bfloat16 |
Gradient norm | 5.0 |
Model | CIDEr | BLEU-4 | METEOR | ROUGE |
Scan2Cap [14] | 35.2 | 22.4 | 21.4 | 43.5 |
3DJCG [6] | 49.5 | 31.0 | 24.2 | 50.8 |
3D-VisTA [117] | 66.9 | 34.0 | 27.1 | 54.3 |
Vote2Cap-DETR++ [12] | 74.4 | 37.2 | 26.2 | 53.3 |
3DVLP [103] | 54.4 | 34.1 | 34.3 | 54.3 |
LEO [34] | 72.4 | 38.2 | 27.9 | 58.1 |
LL3DA [11] | 65.2 | 36.8 | 26.0 | 55.1 |
PQ3D [118] | 80.3 | 36.0 | 29.1 | 57.9 |
Chat-Scene [32] | 77.2 | 36.3 | 28.0 | 58.1 |
LLaVA-3D [115] | 79.2 | 41.1 | 30.2 | 63.4 |
3D-LLaVA [21] | 78.8 | 36.9 | 27.1 | 57.7 |
Video-3D LLM [111] | 80.0 | 40.2 | 28.5 | 61.7 |
GPT4Scene [71] | 86.3 | 40.6 | 28.2 | 59.3 |
LEO-VL | 77.3 | 43.8 | 30.3 | 64.9 |
ScanQA and SQA3D.
We present extended results with more metrics for the experiments reported in the main paper. We report the full results for ScanQA in Appendix C and SQA3D in Appendix C.
Model | CIDEr | BLEU-4 | METEOR | ROUGE | EM | EM-R |
Domain scaling | ||||||
LEO-VL (ScanNet) | 101.1 | 16.2 | 19.6 | 47.4 | 23.0 | 45.9 |
LEO-VL (Scan+3R) | 97.2 | 15.7 | 19.0 | 45.7 | 21.9 | 44.3 |
LEO-VL | 97.6 | 14.8 | 19.2 | 46.1 | 22.2 | 45.0 |
Voxel vs. CFG | ||||||
voxel-full | 101.5 | 15.7 | 20.0 | 48.0 | 23.3 | 46.4 |
voxel-sample | 96.5 | 15.2 | 19.2 | 45.9 | 21.6 | 43.5 |
LEO-VL (ScanNet) | 101.1 | 16.2 | 19.6 | 47.4 | 23.0 | 45.9 |
Position embedding | ||||||
w/o xyz-pos | 97.7 | 14.3 | 19.2 | 46.1 | 21.9 | 44.2 |
w/o z-pos | 97.1 | 14.6 | 19.1 | 46.1 | 21.8 | 44.0 |
LEO-VL (ScanNet) | 101.1 | 16.2 | 19.6 | 47.4 | 23.0 | 45.9 |
Task effects | ||||||
w/o cap | 97.9 | 14.8 | 19.2 | 46.2 | 22.3 | 44.3 |
w/o pl+dl | 95.8 | 13.9 | 18.7 | 45.8 | 21.9 | 44.7 |
LEO-VL (ScanNet) | 101.1 | 16.2 | 19.6 | 47.4 | 23.0 | 45.9 |
Quality vs. scale | ||||||
LEO-VL (Scan+3R) | 97.2 | 15.7 | 19.0 | 45.7 | 21.9 | 44.3 |
+QA | 96.4 | 14.9 | 18.9 | 45.4 | 22.0 | 43.1 |
Data scaling | ||||||
12.5% | 79.2 | 10.6 | 15.9 | 39.6 | 19.4 | 38.3 |
25% | 86.4 | 11.4 | 17.3 | 42.6 | 20.8 | 41.1 |
50% | 94.8 | 13.6 | 18.7 | 44.7 | 21.2 | 42.0 |
LEO-VL | 97.6 | 14.8 | 19.2 | 46.1 | 22.2 | 45.0 |
Post-training on SQA3D | ||||||
SFT | 91.3 | 12.5 | 17.9 | 44.4 | 22.4 | 44.8 |
SceneDPO | 92.8 | 13.1 | 18.4 | 45.1 | 22.7 | 44.4 |
Model | What | Is | How | Can | Which | Others | EM | EM-R |
Domain scaling | ||||||||
LEO-VL (ScanNet) | 52.0 | 71.9 | 55.1 | 67.8 | 54.7 | 57.1 | 58.7 | 61.8 |
LEO-VL (Scan+3R) | 50.2 | 71.8 | 54.0 | 69.2 | 57.8 | 56.5 | 58.3 | 61.3 |
LEO-VL | 52.1 | 72.5 | 56.8 | 70.7 | 59.8 | 56.2 | 59.7 | 62.6 |
Voxel vs. CFG | ||||||||
voxel-full | 51.8 | 71.0 | 55.5 | 66.9 | 49.6 | 58.1 | 58.1 | 61.1 |
voxel-sample | 49.0 | 69.8 | 51.4 | 68.0 | 55.8 | 54.6 | 56.6 | 59.7 |
LEO-VL (ScanNet) | 52.0 | 71.9 | 55.1 | 67.8 | 54.7 | 57.1 | 58.7 | 61.8 |
Position embedding | ||||||||
w/o xyz-pos | 50.5 | 71.6 | 54.2 | 67.2 | 47.3 | 55.5 | 57.0 | 59.7 |
w/o z-pos | 50.9 | 71.6 | 56.6 | 66.3 | 53.8 | 58.0 | 58.4 | 61.2 |
LEO-VL (ScanNet) | 52.0 | 71.9 | 55.1 | 67.8 | 54.7 | 57.1 | 58.7 | 61.8 |
Task effects | ||||||||
w/o cap | 50.1 | 70.9 | 54.4 | 68.9 | 46.4 | 55.3 | 56.8 | 59.8 |
w/o pl+dl | 52.7 | 71.9 | 53.5 | 68.9 | 53.0 | 55.3 | 58.4 | 61.4 |
LEO-VL (ScanNet) | 52.0 | 71.9 | 55.1 | 67.8 | 54.7 | 57.1 | 58.7 | 61.8 |
Quality vs. scale | ||||||||
LEO-VL (Scan+3R) | 50.2 | 71.8 | 54.0 | 69.2 | 57.8 | 56.5 | 58.3 | 61.3 |
+QA | 49.4 | 73.2 | 52.5 | 66.3 | 49.6 | 54.8 | 56.7 | 59.8 |
Data scaling | ||||||||
12.5% | 44.8 | 67.3 | 46.5 | 66.9 | 59.5 | 53.7 | 54.2 | 56.9 |
25% | 46.4 | 67.3 | 48.6 | 64.8 | 53.3 | 55.1 | 54.4 | 57.1 |
50% | 50.2 | 70.1 | 52.7 | 66.9 | 50.7 | 54.6 | 56.6 | 59.3 |
LEO-VL | 52.1 | 72.5 | 56.8 | 70.7 | 59.8 | 56.2 | 59.7 | 62.6 |
Post-training on SQA3D | ||||||||
SFT | 53.0 | 73.3 | 55.5 | 68.3 | 65.2 | 60.8 | 61.0 | 64.0 |
SceneDPO | 53.1 | 73.5 | 54.4 | 68.0 | 66.4 | 61.3 | 61.1 | 64.2 |
MSQA.
We extend the baseline coverage on the MSQA benchmark. Existing baselines for MSQA are relatively scarce, primarily because models generally require training on MSQA to attain reasonable performance, i.e., zero-shot results tend to be poor. As only a handful of recent models have been trained on MSQA, the available baselines are limited. In Appendix C, we report additional baselines and present more detailed metrics on the MSQA ScanNet subset, including GPT-4o [60], LEO [34], MSR3D [46], and SplatTalk [83]. Note that existing evaluation results on MSQA include two different versions of test data: the initially released version and a subsequent refined version. Results corresponding to the earlier version are marked with †.
Model | Count. | Exist. | Attr. | Spatial | Navi. | Others | Overall |
GPT-4o∗‡ [60] | 32.3 | 79.3 | 79.0 | 37.0 | 31.7 | 91.6 | 52.3 |
LEO [34] | 32.5 | 88.5 | 58.7 | 44.2 | 39.6 | 81.4 | 54.8 |
MSR3D [46] | 32.3 | 93.1 | 50.0 | 46.5 | 54.1 | 75.6 | 54.2 |
SplatTalk†‡ [83] | 19.6 | 60.3 | 44.0 | 35.8 | 35.5 | 61.8 | 41.8 |
LEO-VL (ScanNet) | 32.0 | 92.7 | 59.2 | 57.3 | 48.1 | 83.5 | 58.4 |
LEO-VL | 36.3 | 93.8 | 59.2 | 56.0 | 44.6 | 83.5 | 58.2 |
Beacon3D.
Similar to MSQA, the Beacon3D benchmark also has evaluation results based on two versions of test data: an initial version corresponding to the results reported in the paper [35], and a refined version that is released. We denote results corresponding to the earlier version with †. In Appendix C, we report the results on Beacon3D with more detailed metrics.
Model | Class | App. | Geo. | Spa. | Exi. | Overall (case) | Overall (obj.) |
3D-VisTA† [117] | 20.5 | 33.5 | 52.1 | 33.8 | 36.5 | 35.3 | 8.1 |
PQ3D [118] | 37.8 | 45.8 | 32.1 | 19.2 | 44.5 | 35.9 | 4.2 |
SceneVerse [39] | 26.4 | 40.4 | 40.0 | 35.0 | 54.1 | 40.5 | 4.7 |
LEO [34] | 16.4 | 39.8 | 47.6 | 52.8 | 54.3 | 45.2 | 7.5 |
Chat-Scene† [32] | 36.4 | 39.8 | 56.7 | 47.6 | 48.8 | 45.8 | 7.8 |
LEO-VL (ScanNet) | 43.0 | 65.8 | 56.7 | 56.6 | 56.3 | 58.0 | 17.7 |
LEO-VL | 43.0 | 68.6 | 57.8 | 57.2 | 58.0 | 59.5 | 18.2 |
Appendix D Position Embedding
D.1 Rotary Position Embedding
Introduction.
[…]
Formulation.
[…]
where the rotation matrix is block-diagonal, composed of 2D rotation submatrices:
with channel-specific frequencies .
Height Encoding in 3D Scenes.
We adopt RoPE to encode the voxel heights in 3D scenes considering its advantage in capturing height information during voxel feature pooling. As illustrated in Fig. 9, when two objects swap their heights, RoPE-encoded features can distinguish the configurations, whereas additive position embeddings fail. This property motivates our adoption of RoPE for the encoding of height information.
D.2 Fourier Features
Introduction.
Fourier features [80, 44] focus on encoding multi-dimensional spatial positions, such as pixel positions in an image. Classical methods typically process each dimension independently and then combine via concatenation. However, this falls short in capturing holistic spatial relations such as Euclidean distances, and struggles with generalizing to unbounded ranges of multi-dimensional coordinates. Fourier features overcome these issues by mapping the low-dimensional coordinates into a high-dimensional periodic function space, enabling better modeling of multi-dimensional positions.
Formulation.
Given an -dim spatial coordinate and a projection matrix (where is the feature dimension), the Fourier feature is computed as:
where is initialized from a Gaussian distribution. The resulting feature vector is then scaled by for normalization and passed through an MLP to generate the final position embedding, which is subsequently added to the original feature vector .
Appendix E Data Statistics
Additional Simplistic QA Data.
In the main paper, we explore the outcome of naively scaling with low-quality 3D-VL data. The additional QA data is sourced from LEO [34] (94k from 3RScan) and MMScan [54] (575k from ScanNet and 3RScan). This part of data exhibits a simplistic distribution with limited answer diversity. By categorizing answers into placeholder-based templates, we observe a significantly higher proportion of the top-15 frequent answer templates in the additional QA data (39.2% vs. 12.8%). The templates and their distribution are visualized in Figs. 10 and 11.
Appendix F Qualitative Examples
We present qualitative results of LEO-VL performing various tasks in Fig. 12.